# Detecting Viral Social Events through Censored Observation with Deep Survival Analysis

### v1

@@ Abstract
Users increasing activity across various social networks made it the most widely used platform for exchanging and propagating information among individuals. To spread information within a network, a user initially shared information on a social network, and then other users in direct contact with him might have shared that information. Information expanded throughout the network by repeatedly following this process. A set of information that became popular and was repeatedly shared by different individuals was called viral events. Identifying and analyzing viral social events led to valuable insights into the dynamics of information dissemination within a network. However, more importantly, proactive approaches emerged. In other words, by observing the dissemination pattern of a piece of information in the early stages of expansion, it became possible to determine whether this cascade would become viral in the future. This research aimed to predict and detect viral events in social networks by observing granular information and using a deep survival analysis-based method. This model could play a significant role in identifying rumors, predicting the impact of information, and assisting in optimal decision-making in information management and marketing. Ultimately, the proposed method was tested on various real-world datasets from Twitter, Weibo, and Digg.
@@ 1 Introduction
  • Early detection of rumors: Rumors are information with speculative and unverified characteristics. A rumor may be completely false, partially true, or completely accurate in terms of accuracy. When a rumor spreads among users, one of the most critical questions is whether this rumor is a viral event or whether it is merely a non-viral rumor. If a rumor does not infect many users, it is unnecessary to invest time and money to prevent its spread. However, if we know that many users will be exposed to this rumor, it is necessary to manage the crisis and take steps to eliminate its effects or prevent its spread. In cases where this rumor is false information or fake news, the issue becomes more sensitive. We must adopt mitigation approaches if we understand that it will likely become viral.
@@
  • Viral disease prevalence: When a virus spreads through a population, an estimate of virality can be made based on the number of infections at different times. Thus, it can be determined whether there is a need to quarantine or restrict drug production shortly.
  • Traffic congestion: Transport networks include communication links between spatial nodes. Observations of commuting at different times can be used to predict whether traffic and crowding will exceed a defined controllable limit in future hours. Control policies must be applied before congestion occurs if the crowd is expected to exceed the predefined limit.
@@
    Censored data: We can determine the virality of a cascade of events by counting the number of infections if the cascade lifetime is completely observed. However, the main challenge arises when only part of the data is available. This makes monitoring the limited early stages of the cascade life cycle necessary to detect the viral cascades. Studies have shown that viral and non-viral cascades may have the same pattern in the early stages, making it difficult to distinguish between various types of cascades Gao et al. (2022).
@@ 2 Definitions and Preliminary Concepts
Viral Events: When user $u$ disseminates information $i$ over the network at a certain time $t$, a social Event is generated, represented by $(u,t)$. The sequence of social events associated with a particular piece of information $i$ over a network is called a Social Event Cascade. The cascade $c_{i}$ about information ${i}$ with $n$ number of user infections is defined as $c_{i}=\{(u_{0},t_{0}),(u_{1},t_{1}),\dots,(u_{n-1},t_{n-1}),(u_{n},t_{n})\}$ where $t_{0}\leq t_{1}\leq\dots\leq t_{n}$ and $u_{0}$ is the initiator (source) of the cascade. Several of these initiator events are more likely to receive greater attention and reach a wider audience. The cascade of events, in this case, is known as the “viral” phenomenon. Formally, a Viral Social Event is an event $(u_{0},t_{0})$ that initiates a cascade $c_{i}$ in which the amount of infected users ($n$) exceeds a certain threshold of $\zeta$.
@@
[…]
@@ 4 Proposed Method
Figure 2: A brief overview of the VEDSA method. A cascade of inputs is observed between $[0,\tau]$, corresponding to steps 00 to $r$ after prepossessing by Figure 1. From the early stage of uncensored data, $\gamma$ fits the survival function and predicts the probability of virality for censored data. A discriminator $\delta$ will learn to output the virality label by classifying the cascades based on the estimated survival functions.
@@ 4.1 $\gamma$ Model: Survival Function Fitting
To fit a good survival function, we have three choices Xu & Guo (2023); Moreno (2018):
@@
    Non-parametric: A function is learned without assumptions about distributions and without incorporating covariates, making it more complex as the number of observations increases. As a result, they have greater flexibility. A fitted non-parametric survival function may be a piecewise constant in the interval without observation. In this case, the survival function is called a non-smooth survival function, leading to less stable parameter estimation and making it more challenging to compare different survival functions. In this category, Kaplan-Meier Kaplan & Meier (1958) is one of the best-known methods.
@@
$\sigma^{(i)}_{j}=\begin{cases}1&\text{cascade is in viral state at time $j\geq t^{(i)}_{v}$}\\ 0&\text{\parbox[t]{151.76964pt}{cascade does not reach viral state till time $j<t^{(i)}_{v}$ or the cascade is non-viral $n<\zeta$}}\end{cases}$ (12)

### v2

# Detecting Popular Social Events through Limited Observation with Deep Survival Analysis

@@ Abstract
Users increasing activity across various social networks made it the most widely used platform for exchanging and propagating information among individuals. To spread information within a network, a user initially shared information on a social network, and then other users in direct contact with him might have shared that information. Information expanded throughout the network by repeatedly following this process. A set of information that became popular and was repeatedly shared by different individuals was called popular trends. Identifying and analyzing these trends led to valuable insights into the dynamics of information dissemination within a network. However, more importantly, proactive approaches emerged. In other words, by observing the dissemination pattern of a piece of information in the early stages of expansion, it became possible to determine whether this cascade would become highly popular in the future. This research aimed to predict and detect popular trends in social networks by observing limited early-stage data and using a deep survival analysis-based method. This model could play a significant role in improving recommendation systems, predicting the reach of digital content, and assisting in optimal decision-making in digital marketing. Ultimately, the proposed method was tested on various real-world anonymized datasets from Twitter, Weibo, and Digg.
@@ 1 Introduction
  • Traffic congestion: Transport networks include communication links between spatial nodes. Observations of commuting at different times can be used to predict whether traffic and crowding will exceed a defined controllable limit in future hours. Control policies for traffic management must be applied before congestion occurs if the crowd is expected to exceed the predefined limit.
@@
    Limited data: We can determine the virality of a cascade of events by counting the number of infections if the cascade lifetime is completely observed. However, the main challenge arises when only part of the data is available. This makes monitoring the limited early stages of the cascade life cycle necessary to detect the viral cascades. Studies have shown that viral and non-viral cascades may have the same pattern in the early stages, making it difficult to distinguish between various types of cascades Gao et al. (2022).
@@ 2 Definitions and Preliminary Concepts
Viral Events: When user $u$ disseminates information $i$ over the network at a certain time $t$, a social Event is generated, represented by $(u,t)$. The sequence of social events associated with a particular piece of information $i$ over a network is called a Social Event Cascade. The cascade $c_{i}$ about information i{i} with $n$ number of user infections is defined as $c_{i}=\{(u_{0},t_{0}),(u_{1},t_{1}),\dots,(u_{n-1},t_{n-1}),(u_{n},t_{n})\}$ where $t_{0}\leq t_{1}\leq\dots\leq t_{n}$ and $u_{0}$ is the initiator (source) of the cascade. Several of these initiator events are more likely to receive greater attention and reach a wider audience. The cascade of events, in this case, is known as the “viral” phenomenon. Formally, a Viral Social Event is an event $(u_{0},t_{0})$ that initiates a cascade $c_{i}$ in which the amount of infected users ($n$) exceeds a certain threshold of $\zeta$.
@@
$\begin{aligned} \frac{P(t\leq T<t+dt|T\geq t)}{dt}&=\frac{P(t\leq T<t+dt,t\leq T)}{P(T\geq t)dt}\\ &=\frac{P(t\leq T<t+dt)}{S(t)dt}\end{aligned}$ (5)
@@ 4 Proposed Method
Figure 2: A brief overview of the VEDSA method. A cascade of inputs is observed between $0,\tau]$, corresponding to steps 0 to $r$ after prepossessing by Figure 1. From the early stage of uncensored data, $\gamma$ fits the survival function and predicts the probability of virality for censored data. A discriminator $\delta$ will learn to output the virality label by classifying the cascades based on the estimated survival functions.
@@ 4.1 $\gamma$ Model: Survival Function Fitting
To fit a good survival function, we have three choices Xu and Guo (2023); Moreno (2018):
@@
    Non-parametric: A function is learned without assumptions about distributions and without incorporating covariates, making it more complex as the number of observations increases. As a result, they have greater flexibility. A fitted non-parametric survival function may be a piecewise constant in the interval without observation. In this case, the survival function is called a non-smooth survival function, leading to less stable parameter estimation and making it more challenging to compare different survival functions. In this category, Kaplan-Meier Kaplan and Meier (1958) is one of the best-known methods.
@@
[…]